A rope over the top of a fence has the same length on each side and weighs one third of a pound per foot. On one end hangs a monkey holding a banana, and on the other end a weight equal to the weight of the monkey. The banana weighs 2 ounces per inch. The length of the rope in feet is the same as the age of the monkey ,and the weight of the monkey in ounces is as much as the age of the monkey's mother. The combined ages of the monkey and it's mother are 30 years. One-half the weight of the monkey plus the weight of the banana is one-forth the sum of the weights of the rope and the weight.The monkey's mother is one half as old as the monkey will be when it is three times as old as it's mother was when she was one half as old as the monkey will be when it is as old as its mother will be when she is four times as old as the monkey was when it was twice as old as its mother was when she was one-third as old as the monkiey was when it was as old as its mother was when she was three times as old as the monkey was when it was one-fourth as old as it is now.

HOW LONG IS THE BANANA?

Found 2 solutions by wuwei96815, monkeybusiness:
Answer by wuwei96815(245)

If there is more weight on one side of the wall than the other, why doesn't the monkey fall? Are we suppose to assume that friction is compensating for the weight of the banana?

Answer by monkeybusiness(1)

Alrighty, #1 you can solve this if you use your brain! #2 The banana will not cause it to fall because of the friction on the fence. (it could be a barbed wire fence by the way)
PART ONE
Let x = mother's age now. Let y = offspring's age now
:
The monkey’s mother is half as old as the monkey will be when it is three times as old as its mother was when she was half as old as the monkey will be when it is twice as old as it is now.
:
The last phrase:"when she was half as old as the monkey will be when it is twice as old as it is now." just means: .5(2y) = y
:
Replacing that phrase with y:
mother is half as old as the monkey will be when it is three times as old as its mother was y
:
x = .5(3y)
x = 1.5y
:
"The combined ages of the monkey and its mother are thirty years."
x + y = 30

1.5y + y = 30; substituted 1.5y for x
2.5y = 30
y = 30/2.5
y = 12 yrs old
Then
x = 18 yrs old
PART TWO
I think you should be able to do the rest (its easier from here), don't listen to those people who say it is not possible.